Sndi Inc. has following information
Sales Revenue $2,000,000
Cost of goods sold $1,100,000
Operating Expenses $800,000
Average Capital Invested $1,000,000

1. Compute the Firm's sales margin, Capital turnover and ROI?
2. If the sales and average Invested Capital remains the same Next year, to what level would the total expenses have to be changed so as to improve the Firm's ROI to 15%?
1
Expert's answer
2019-12-26T14:07:07-0500

Answer :

Question 01

Operating income = Sales revenue - (Cost of goods sold + Operating expenses)

Therefore,

Operating income ($) = "2,000,000 - (1,100,000 + 800,000) = 100,000"

Sales margin = (Operating income) / (Sales revenue)

Therefore,

Sales margin = "\frac{100,000}{2,000,000} = 0.05"

Sales margin = 5%


Capital turnover = (Sales revenue) / (Average capital invested)

Capital turnover = "\frac{2,000,000}{1,000,000} = 2"

Capital turnover = 2


ROI = (Operating income) / (Average capital invested)

ROI = "\frac{100,000}{1,000,000} = 0.1"

ROI = 10%


Question 02

New ROI = 15% and the average capital invested will remain the same.

Therefore, the new operating income would be

New operating income = (ROI) * (Average capital invested)

New operating income ($) = "0.15 * 1,000,000 = 150,000"

Since the total sales will also remain constant,

New total expenses would be,

Total expenses ($) = Sales revenue - Operating income

Total expenses ($) = "2,000,000 - 150,000 = 1,850,000"

Total expenses = $ 1,850,000